# A-IPO: Adaptive Intent-driven Preference Optimization

### v1

@@ 1 Introduction
Large language models (LLMs) have seen rapid adoption in fields such as natural language processing, healthcare, finance etc., where they excel at generating human-like text, automating complex tasks, and supporting decision-making (Brown et al., 2020; Bommasani et al., 2021; OpenAI, 2023; Yang et al., 2024b). However, effectively deploying LLMs across diverse domains and regions remains challenging. Key obstacles include domain-specific requirements, linguistic and cultural differences, ethical and safety concerns, and varying regulatory standards (Yang et al., 2024a; Ali et al., 2024; Cheng et al., 2024). These challenges remain unresolved in current research and practice (Bommasani et al., 2021; Weidinger et al., 2021; Liang et al., 2022). Recently, there have been many attempts for domain adaptation (Gururangan et al., 2020; Lee et al., 2020), cultural and linguistic localization (Ahuja, 2023; Chen et al., 2022), mitigation of ethical biases and safety risks (Gehman et al., 2020; Hendrycks et al., 2021), and robustness evaluation (Ribeiro et al., 2020; Zou et al., 2023). Nonetheless, the technology has not yet achieved a point where it can comprehensively address the full range of challenges, making this an ongoing open research problem.
A widely used strategy for aligning LLMs is to incorporate human feedback through preference optimization (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022). For this, Direct Preference Optimization (DPO) has emerged as a de-facto standard method in this area, as it efficiently tunes model outputs to better match human values and judgments (Rafailov et al., 2023; Ziegler et al., 2019). By leveraging explicit pairwise comparisons or ranked preferences, DPO directly adjusts model parameters to align responses with human evaluators’ choices (Ouyang et al., 2022; Rafailov et al., 2023). This enables rapid, targeted improvements in model safety, helpfulness, and alignment with user expectations (Askell et al., 2021; Bai et al., 2022). Below, we summarize the key limitations of DPO and its variants (Rafailov et al., 2023; Yao et al., 2025); a formal and detailed discussion can be found in Appendix B.
@@ 2 Related Work
Preference Alignment. As LLMs become more widely used, aligning their outputs with human values and intentions is critical for safety and trust. Early methods used supervised fine-tuning with human-labeled data (Ouyang et al., 2022; Wei & et al., 2021), but struggled to capture the full range of human preferences. Reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020) improved alignment by optimizing models based on human preference comparisons. Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its variants further streamlined this process by directly optimizing model outputs using pairwise preference data, removing the need for explicit reward modeling and improving alignment with complex values. However, standard DPO assumes uniform preferences, often favoring majority views and neglecting minority or outlier needs (Liu & et al., 2021), raising fairness concerns. To address this, group-based and distributional methods have been developed to explicitly balance performance across user groups.
Group/Pluralistic Alignment. Pluralistic preference optimization approaches were introduced to ensure that language models can accommodate and respect the diverse and sometimes conflicting preferences present within user populations. Chronologically, early work such as EM-DPO (Chidambaram et al., 2024) broke DPO’s homogeneity assumption by learning distributions of different preference types and corresponding response strategies. Building on this, Minmax-DPO (Chidambaram et al., 2024) adopted a “minimax cross-subgroup regret” strategy to balance preference expressions across groups, especially minorities, though both methods faced challenges in computational complexity and subgroup definition dependency. To the best of our knowledge, the most recent advancement in this line of work is GDPO (Yao et al., 2025), which proposes a principled two-stage approach: belief distribution prediction and belief-conditioned response generation, to ensure that minority preferences are adequately addressed. However, its performance heavily relies on accurate belief/group partitioning; moreover, GDPO does not incorporate belief/group information directly into the reward modeling process, which limits its ability to fully leverage group distinctions for preference optimization.
@@ 4.2 Intention module
Here, LLM is prompted with $P_{\text{decomp}}$ to decompose $x$ into sub-questions ($x_{\text{aug}}$), enabling a more precise understanding of user intent (see Figure 1). We also retrieve relevant external information $x_{\text{ext}}$ from Wikipedia knowledge base (Wikimedia Foundation, 2025; Ali et al., 2020), indexed via Pinecone (Pinecone Systems, Inc., 2023)) to provide supporting evidence. The combined input, $x_{con}=\texttt{concat}(x_{\text{aug}},x_{\text{ext}})$, merges the decomposed prompt and retrieved context. Sentence-level fact-checking is applied to $x_{con}$, retaining only statements verified as true via Anah-v2 (Gu et al., 2024). This fact-checked content is passed to the intention module, ensuring intention modeling is based on accurate, reliable evidence. The specific prompt template and generation protocol are detailed in Appendix I.
@@ 7.1 Experimental Settings
(2) Evaluation Metrics. For performance evaluation of A-IPO, we use the following metrics: (i) Win Rate (Dudík et al., 2015), (ii) Intention-Consistency Score (ICS) (Yao et al., 2025), (iii) Response-Intention Consistency (RIC), (iv) Response Similarity(RS) (Yao et al., 2025), (v) Defense Success Rate (DSR) (Wang et al., 2024). Detailed description and mathematical formulations of these metrics are provided in Appendix G.1.2.
(3) Baselines. For performance comparison, we use the following methods as baselines: (i) DPO (Rafailov et al., 2023), (ii) GDPO (Yao et al., 2025), (iii) Few-shot Prompts, (iv) SFT. Detailed description about these baselines is provided in Appendix G.1.3.
@@ Appendix A Background
In this section, we provide a quick background of the widely used preference alignment pipeline: RLHF (Ouyang et al., 2022) as well as the widely used preference alignment methods: Direct Preference Optimization (DPO) (Rafailov et al., 2023) and GDPO (Yao et al., 2025).
@@ A.3 Group Distributional Preference Optimization (GDPO)
GDPO (Yao et al., 2025) addresses DPO’s limitations by explicitly modeling the diversity of human preferences. Instead of directly learning $p_{\theta}(y|x)$, GDPO decomposes response generation into two steps: (1) predicting a belief distribution over preference categories $p_{\theta}(b|x)$, and (2) generating a response conditioned on both the input and a sampled belief $p_{\theta}(y|b,x)$. The overall response distribution is thus a mixture:
@@ G.1.2 Evaluation Metrics
(iv) Response Similarity (RS) (Yao et al., 2025): RS evaluates the semantic similarity between the model’s response $\hat{y}_{i}$ and a reference response $y_{i}$ (both expressing the same intention $\mathcal{I}$). We use Sentence-BERT (all-mpnet-base-v2) to obtain embeddings, tokenize each response (up to 512 tokens), extract the [CLS] embedding, and apply L2 normalization. The cosine similarity is computed as:
@@ G.1.3 Baselines
(i) GDPO (Yao et al., 2025). GDPO extends DPO by explicitly modeling group-level belief distributions. It employs a two-stage process: first calibrating belief predictions, then aligning responses conditioned on these beliefs. This baseline enables a direct comparison with A-IPO’s approach to capturing implicit group or community preferences, especially in scenarios where group labels are not pre-defined.
@@ G.1.4 Experimental Setup
Policy and Reference Models. Similar to DPO (Rafailov et al., 2023) and GDPO (Yao et al., 2025), we use GPT2-Large (Radford et al., 2019) (774M parameters) and Pythia-2.8B (Biderman et al., 2023), as our target LLMs. The reference model is obtained by supervised fine-tuning on $(x,y_{w})$.

### v3

@@ 1 Introduction
Large language models (LLMs) have seen rapid adoption in fields such as natural language processing, healthcare, finance etc., where they excel at generating human-like text, automating complex tasks, and supporting decision-making (Brown et al., 2020; Achiam et al., 2023). However, effectively deploying LLMs across diverse domains and regions remains challenging. Key obstacles include domain-specific requirements, linguistic and cultural differences, ethical and safety concerns, and varying regulatory standards. These challenges remain unresolved in current research and practice (Weidinger et al., 2021; Ali et al., 2025; Liang et al., 2022). Recently, there have been many attempts for domain adaptation (Gururangan et al., 2020; Lee et al., 2020), cultural and linguistic localization (Ahuja et al., 2023; Wang et al., 2019), mitigation of ethical biases and safety risks (Gehman et al., 2020; Hendrycks et al., 2021), and robustness evaluation (Ribeiro et al., 2020; Zou et al., 2023). Nonetheless, the technology has not yet achieved a point where it can comprehensively address the full range of challenges, making this an ongoing open research problem.
A widely used strategy for aligning LLMs is to incorporate human feedback through preference optimization (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022). For this, Direct Preference Optimization (DPO) has emerged as a de-facto standard method in this area, as it efficiently tunes model outputs to better match human values and judgments (Rafailov et al., 2023; Ziegler et al., 2019). By leveraging explicit pairwise comparisons or ranked preferences, DPO directly adjusts model parameters to align responses with human evaluators’ choices (Ouyang et al., 2022; Rafailov et al., 2023). This enables rapid, targeted improvements in model safety, helpfulness, and alignment with user expectations (Askell et al., 2021; Bai et al., 2022). Below, we summarize the key limitations of DPO and its variants (Rafailov et al., 2023; Yao et al., 2024); a formal and detailed discussion can be found in Appendix B.
@@ 2 Related Work
Preference Alignment. As LLMs become more widely used, aligning their outputs with human values and intentions is critical for safety and trust. Early methods used supervised fine-tuning with human-labeled data (Ouyang et al., 2022; Wei et al., 2021), but struggled to capture the full range of human preferences. Reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020) improved alignment by optimizing models based on human preference comparisons. Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its variants further streamlined this process by directly optimizing model outputs using pairwise preference data, removing the need for explicit reward modeling and improving alignment with complex values. However, standard DPO assumes uniform preferences, often favoring majority views and neglecting minority or outlier needs, raising fairness concerns. To address this, group-based and distributional methods have been developed to explicitly balance performance across user groups.
Group/Pluralistic Alignment. Pluralistic preference optimization approaches were introduced to ensure that language models can accommodate and respect the diverse and sometimes conflicting preferences present within user populations. Chronologically, early work such as EM-DPO (Chidambaram et al., 2024) broke DPO’s homogeneity assumption by learning distributions of different preference types and corresponding response strategies. Building on this, Minmax-DPO (Chidambaram et al., 2024) adopted a “minimax cross-subgroup regret” strategy to balance preference expressions across groups, especially minorities, though both methods faced challenges in computational complexity and subgroup definition dependency. To the best of our knowledge, the most recent advancement in this line of work is GDPO (Yao et al., 2024), which proposes a principled two-stage approach: belief distribution prediction and belief-conditioned response generation, to ensure that minority preferences are adequately addressed. However, its performance heavily relies on accurate belief/group partitioning; moreover, GDPO does not incorporate belief/group information directly into the reward modeling process, which limits its ability to fully leverage group distinctions for preference optimization.
@@ 4.2 Intention module
Here, LLM is prompted with $P_{\text{decomp}}$ to decompose $x$ into sub-questions ($x_{\text{aug}}$), enabling a more precise understanding of user intent (see Figure 1). We also retrieve relevant external information $x_{\text{ext}}$ from Wikipedia [^1] https://dumps.wikimedia.org/ knowledge base (Ali et al., 2020), indexed via Pinecone [^2] https://www.pinecone.io/) to provide supporting evidence. The combined input, $x_{con}=\texttt{concat}(x_{\text{aug}},x_{\text{ext}})$, merges the decomposed prompt and retrieved context. Sentence-level fact-checking is applied to $x_{con}$, retaining only statements verified as true via Anah-v2 (Gu et al., 2024). This fact-checked content is passed to the intention module, ensuring intention modeling is based on accurate, reliable evidence. The specific prompt template and generation protocol are detailed in Appendix I.
@@ 7.1 Experimental Settings
(2) Evaluation Metrics. For performance evaluation of A-IPO, we use the following metrics: (i) Win Rate (Dudík et al., 2015), (ii) Intention-Consistency Score (ICS) (Yao et al., 2024), (iii) Response-Intention Consistency (RIC), (iv) Response Similarity(RS) (Yao et al., 2024), (v) Defense Success Rate (DSR) (Wang et al., 2024). Detailed description and mathematical formulations of these metrics are provided in Appendix G.1.2.
(3) Baselines. For performance comparison, we use the following methods as baselines: (i) DPO (Rafailov et al., 2023), (ii) GDPO (Yao et al., 2024), (iii) Few-shot Prompts, (iv) SFT. Detailed description about these baselines is provided in Appendix G.1.3.
@@ Appendix A Background
In this section, we provide a quick background of the widely used preference alignment pipeline: RLHF (Ouyang et al., 2022) as well as the widely used preference alignment methods: Direct Preference Optimization (DPO) (Rafailov et al., 2023) and GDPO (Yao et al., 2024).
@@ A.3 Group Distributional Preference Optimization (GDPO)
GDPO (Yao et al., 2024) addresses DPO’s limitations by explicitly modeling the diversity of human preferences. Instead of directly learning $p_{\theta}(y|x)$, GDPO decomposes response generation into two steps: (1) predicting a belief distribution over preference categories $p_{\theta}(b|x)$, and (2) generating a response conditioned on both the input and a sampled belief $p_{\theta}(y|b,x)$. The overall response distribution is thus a mixture:
@@ G.1.2 Evaluation Metrics
(iv) Response Similarity (RS) (Yao et al., 2024): RS evaluates the semantic similarity between the model’s response $\hat{y}_{i}$ and a reference response $y_{i}$ (both expressing the same intention $\mathcal{I}$). We use Sentence-BERT (all-mpnet-base-v2) to obtain embeddings, tokenize each response (up to 512 tokens), extract the [CLS] embedding, and apply L2 normalization. The cosine similarity is computed as:
@@ G.1.3 Baselines
(i) GDPO (Yao et al., 2024). GDPO extends DPO by explicitly modeling group-level belief distributions. It employs a two-stage process: first calibrating belief predictions, then aligning responses conditioned on these beliefs. This baseline enables a direct comparison with A-IPO’s approach to capturing implicit group or community preferences, especially in scenarios where group labels are not pre-defined.
@@ G.1.4 Experimental Setup
Policy and Reference Models. Similar to DPO (Rafailov et al., 2023) and GDPO (Yao et al., 2024), we use GPT2-Large (Radford et al., 2019) (774M parameters) and Pythia-2.8B (Biderman et al., 2023), as our target LLMs. The reference model is obtained by supervised fine-tuning on $(x,y_{w})$.